OpenStudy (anonymous):

A student is riding a bike at a speed of 17 mph. How many minutes will it take for the student to travel 2.5 miles to a friends house? @Precalculus Pre-A…

OpenStudy (saifoo.khan):

\[time = \frac{distance}{speed}\]

OpenStudy (saifoo.khan):

\[time = \frac{2.5}{17}\] \[time = \frac{5}{34}hours\]

OpenStudy (saifoo.khan):

in minutes, \[\frac{5}{34} \times 60\] \[\frac{150}{17}=8.82mins\]

OpenStudy (anonymous):

8.7